1.lynne jogs daily around a circular track with a diameter of 25 m for a total distance of 942 m in 5.0 min.


a.How many times did lynne jog around track?


b.what is lynne's average angular speed in radians/s?


1
Expert's answer
2021-12-02T10:08:22-0500
"N=\frac{942}{\pi (25)}=12\\
\omega=\frac{2\pi (12)}{5\cdot60}=0.25\frac{rad}{s}"